OpenStudy (anonymous):

simplify 3(5x+2) i think it is 21x

OpenStudy (anonymous):

i don't

OpenStudy (anonymous):

Me neither.

OpenStudy (anonymous):

then can you explain it to me

OpenStudy (anonymous):

distribute the \(3\)

OpenStudy (anonymous):

\[a(b+c)=ab+ac\] \[3(5x+2)=3\times 5x+3\times 2\]

OpenStudy (anonymous):

so 5x +6

OpenStudy (anonymous):

no